OpenStudy (anonymous):

Solve the system of equations 2x - y + 2z = 15 -x + y + z = 3 3x - y + 2z = 18

OpenStudy (anonymous):

z=3+x-y Substitute 2x - y + 2(3+x-y) = 15 3x - y + 2(3+x-y) = 18 &Solve

OpenStudy (anonymous):

^This